OpenStudy (anonymous):

Mary Has a bag that contains 4 green marbles, 8 red marbles, and 9 blue marbles mary draws a marble from the bag, what is the probability that she will draw a green or blue marble?

OpenStudy (anonymous):

picture of problem.

OpenStudy (anonymous):

number of green or blues divided by total number

OpenStudy (anonymous):

Total balls is 21 P(green+blue)=4/21+9/13=13/21

OpenStudy (anonymous):

thanks